You are on an airplane traveling 30° south of due west at 100 m/s with respect to the air. The air is moving with a speed 36 m/s with respect to the ground due north. 1) What is the speed of the plane with respect to the ground? 87.73 m/s Submit Help 2) What is the heading of the plane with respect to the ground? (Let 0° represent due north, 90 represents due east). 74.5 East of due North Submit | | Help | 3) How far west will the plane travel in 1 hour? m Submit Help
Answer #1

Solution)

Given,

Vp = velocity of plane =36 m/s at angle = 30 degree south due west

Va = velocity of air = 36 m/s

1) Vy = - 100 sin 30 + 36= - 14 m/s

Vx = - 100*cos 30 = -86.6 m/s

V = speed of plane = ( Vx^2 + Vy^2 )^1/2

= 87.72 m/s

C) distance travel in west= Vx*t

= - 86.6*3600sec = - 311760 m (Ans)

[ -ve sign shows in west direction]

=======

Good luck!:)